OpenStudy (anonymous):

Choose the correct simplification of the expression 4 over y to the power of negative 3. 4y3 4 over y to the 3rd power y to the 3rd power over 4 This expression is already simplified.

hero (hero):

\[\frac{4}{y^{-3}}\]

hero (hero):

The negative exponent means "inverse". Also there's a rule of exponents: \[\frac{1}{x^{-a}} = x^a\]

hero (hero):

In other words, the inverse of one over x to the power a equals x to the power a.

OpenStudy (anonymous):

so what is the answer